OpenStudy (anonymous):

Help please.. I want to say B, but that would mean they aren't girls or boys.. Of the 150 students at Smith Elementary School, 22% were absent from school on Tuesday. Of the students who were absent, 20 of them were male. Which claim cannot be made based on the given information? A. There are 150 students at Smith Elementary School. B. 13 females were absent on Tuesday. C. 78% of the students were not absent on Tuesday. D. 78 of the students who were not absent on Tuesday were male.

OpenStudy (mathstudent55):

In choice D you mean 78% right?

OpenStudy (anonymous):

It just says 78 not 78% @mathstudent55 That's what I got confused on as well

OpenStudy (mathstudent55):

Ok. Let's look at each choice and see if we can conclude it from the given info. The one you can't conclude is the answer.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

We are told A is true. 100-22=78 so C is true. 150 * 22% = 33 students absent. 150-33=117 students not absent. 33 absent - 20 boys = 13 girls absent. So B is true. Only one left is D.

OpenStudy (mathstudent55):

Choice A. There are 150 students at Smith Elementary School. The problem starts with "Of the 150 students at Smith Elementary School..." Obviously there are 150 students at the school. Choice A is not it.

OpenStudy (anonymous):

Yep, I got that

OpenStudy (mathstudent55):

Choice B. 13 females were absent on Tuesday. 22% of the 150 students were absent. 22% of 150 = 0.22 * 150 = 33 33 students were absent. Of the 33 absent students, 20 are male. That means 13 absent students are female. We can conclude choice B, so that's not the answer.

OpenStudy (anonymous):

Okay. That makes sense

OpenStudy (mathstudent55):

Choice C. 78% of the students were not absent on Tuesday. If 22% of the students were absent, then 100% - 22% = 78%. That means 78% of the students were not absent. We can conclude choice C, so that is not the answer.

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

By default, the only choice left is D, so D must be the answer. Now let's look at it just to confirm.

OpenStudy (anonymous):

Ok

OpenStudy (mathstudent55):

Choice D. 78 of the students who were not absent on Tuesday were male. 78% of the 150 students were not absent. 78% of 150 = 0.78 * 150 = 117 All we know about the students who were not absent is that there are 117 of them. We are not told what the break down of male and female students is for the students who were not absent. That's why we can't conclude that 78 of the present students are male.

OpenStudy (anonymous):

That makes more sense now

OpenStudy (mathstudent55):

Great.

OpenStudy (anonymous):

Could you help me on some more questions that I don't get.

OpenStudy (mathstudent55):

Sure. Just start a new post. I'll go there.

OpenStudy (anonymous):

Thanks! :)

OpenStudy (mathstudent55):

wlcm